# Exact Learning Dynamics of In-Context Learning in Linear Transformers and Its Application to Non-Linear Transformers

### v1

@@ 1 Introduction
The transformer architecture [1] has revolutionized machine learning, particularly through its capability for in-context learning (ICL). ICL allows models to adapt their behavior based on examples provided within the input context, obviating the need for task-specific parameter updates [2]. This phenomenon links conceptually to meta-learning, characterized by distinct timescales: slow learning via weight updates during training and fast adaptation based on context during inference [3]. Unlike traditional supervised learning, ICL enables generalization to novel tasks by conditioning on contextual examples [4]. Mechanistically, attention layers can implement adaptive processing analogous to fast weights [5, 6, 3], suggesting that transformers might internally simulate learning algorithms [7, 3, 8]. Despite compelling empirical demonstrations, a full theoretical understanding of how transformers achieve ICL remains unclear.
To address this challenge, our work presents an in-depth analysis of a simplified, mathematically tractable transformer model performing in-context linear regression. Our model incorporates specific simplifications—linear attention [3, 7], a structured input embedding [9, 7], mean-squared error loss, and a continuous-time approximation of learning dynamics—yet retains core aspects of ICL. This deliberate simplification allows us to analytically trace how contextual information modulates the model’s predictions. By using methods from the study of exactly solvable models, introduced first in deep linear networks [10], we derive a precise characterization of the learning dynamics under stochastic gradient descent (SGD). We analyze the resulting phenomenology, including the emergence of multiple learning timescales dictated by data statistics [10, 11]. This detailed understanding of a simplified system aims to provide basic insights applicable to the interpretation and analysis of more complex transformer models exhibiting ICL. In particular, we believe that this detailed understanding can inform interpretability efforts and evaluation methodologies for more complex, real-world models and argue the point by quantitatively analyzing the properties of model parameters in 1) simple attention only transformer model that exhibit sudden emergence of in context learning during training time, and 2) grokking model that exhibits delayed generalization.
@@ 1.1 Related Work
Transformers exhibit remarkable in-context learning (ICL), adapting to new tasks using context examples without explicit parameter updates [2]. This connects to classic idea of meta-learning’s dual timescales: slow weight learning and fast context adaptation [3]. Early concepts like fast weights [5] suggested context-based adaptation mechanisms. Recent studies demonstrate that linear self-attention layers can be mathematically equivalent to fast weight update rules [6, 3], enabling Transformers to implement learning algorithms within their forward pass [3].
@@
Several studies confirm that Transformers can encode standard learning algorithms. Akyürek et al. [7] proved by construction that Transformers can implement gradient descent (GD) and ridge regression, finding empirically that trained models mimic these algorithms and converge towards Bayesian optimal predictors for linear tasks. Similarly, Von Oswald et al. [3] demonstrated that Transformers trained on regression problems learn GD internally, with attention updates mapping directly to GD steps, effectively becoming "mesa-optimizers". In parallel work Ahn et al. [8], and Zhang et al. [12] provided theoretical guarantees that standard stochastic gradient descent (SGD) training yields preconditioned GD at the fixed-point of the training dynamics across a range of setups. These findings collectively suggest that standard training procedures can induce Transformers to implicitly learn optimization algorithms. Extending this, Fu et al. [13] found evidence that Transformers might learn higher-order methods resembling Newton’s method, explaining superior performance on ill-conditioned problems. Alternative perspectives conceptualize ICL as Bayesian inference, where models infer latent task variables from prompts [14], or as kernel regression, where attention patterns reflect similarity-based weighting akin to kernel methods [15]. These views provide explanations for observed phenomena like prompt sensitivity and the utility of relevant examples [14, 15]. The precise mechanism—gradient-based, Bayesian, or kernel-like—likely depends on model scale, architecture, and training data [7, 3, 15, 14, 8, 12].
@@
At the same time, analyzing exact learning dynamics has a rich history, notably through the study of deep linear networks [10]. Saxe et al. [10] derived exact solutions for these models, revealing nonlinear dynamics and multiple timescales despite linearity. Learning often occurs in stages, marked by plateaus followed by rapid error reduction as different eigenmodes of the data correlation are sequentially learned. Learning timescales are typically eigenvalue-dependent, with dominant modes learned faster [10, 16]. In nonlinear networks, exact solutions are rare, but phenomena like spectral low rank bias indicate structured learning dynamics: models tend to learn low-frequency (smooth) function components before high-frequency details [11]. Such low rank bias is aslo linked to the Neural Tangent Kernel (NTK) spectrum of highly non-linear models, where larger eigenvalues associated with smoother functions lead to faster learning [11, 17], potentially explaining the generalization capabilities of overparameterized models.
@@
Separation of timescales emerges as a common theme across linear dynamics [10], nonlinear spectral bias [11], meta-learning [3] and the abrupt emergence of ICL or induction heads [18] hints at similar timescale separation. Indeed, there have been theoretical results in non-linear transformer establishing distinct phases in the learning dynamics [19, 20] further supporting the idea. Alternative models for ICL like associative memories instead connect ICL to retrieval dynamics [21].
@@ 2.1 Linear Transformer Architecture
We analyze a single-layer transformer with a single linear attention head. In contrast to standard architectures using softmax attention [1], we employ simplifications common in theoretical studies [3, 7]. The model utilizes two weight matrices, $W^{Q}\in\mathbb{R}^{D\times D}$ (query/key) and $W^{P}\in\mathbb{R}^{D\times D}$ (output/projection), and replaces the attention non-linearity with the identity function ($\phi(x)=x$). The layer’s computation is defined as (see Appendix A.1-A.3, Eq. (A3)):
@@
Here, $Z\in\mathbb{R}^{D\times(N+1)}$ is the input sequence embedding, and $N+1$ is the sequence length. The learnable parameters are $\Theta=\{W^{P},W^{Q}\}$. This linear attention model, while simplified, is sufficient for learning in-context linear regression [3].
Relevant parameter submatrices are defined based on the input embedding structure (Appendix A.2). We partition $W^{Q}$ such that its first $d$ columns form $q=\left[\;q_{1}\,\mid\,q_{2}\;\right]^{T}$, and $W^{P}$ such that its bottom $d$ rows form $p=\left[\;p_{1}\,\mid\,p_{2}\;\right]$. All submatrices $q_{1},q_{2},p_{1},p_{2}$ are in $\mathbb{R}^{d\times d}$. More precisely, we have the following parametrization:
$W^{P}$ =($\ p_{1}&p_{2}\end{pmatrix}\quad\text{with}\quad p=\begin{pmatrix}p_{1}&p_{2}\end{pmatrix},$ (2)
@@ 2.2 Embedding Structure and Task Setup for In-Context Regression
The model is trained to perform in-context linear regression. Given a prompt containing $N$ example pairs $(x_{i},y_{i})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, it must predict the output $y_{q}$ for a query input $x_{q}\in\mathbb{R}^{d}$ based on the linear relationship demonstrated in the prompt. This setup extends the $\mathbb{R}^{d}\to\mathbb{R}^{1}$ tasks studied in some prior work [7, 9] to the more general vector-valued regression setting $\mathbb{R}^{d}\to\mathbb{R}^{d}$.
We employ a structured embedding strategy commonly used in ICL studies [7, 9, 22]. For an input sequence $(x_{1},y_{1},\dots,x_{N},y_{N},x_{q})$, the embedding matrix $Z\in\mathbb{R}^{D\times(N+1)}$ (with $D=2d$) is constructed as:
@@ 3 Theory
Simplified predictor. The model predicts the query output $\hat{y}\in\mathbb{R}^{d}$ using the bottom $d$ elements of the final output column $[f(Z)]_{:,N+1}$. As derived in Appendix A.9 (Eq. (A18)), this prediction simplifies to a linear transformation of the query input:
@@
Null Parameter Dynamics and Predictor Simplification. A key simplification arises from identifying $p_{1}$ and $q_{2}$ as "null parameters." Our analysis (Appendix A.10) demonstrates that if initialized at zero, their expected gradients under SGD vanish [12, 8]. This theoretically justifies fixing $p_{1}=\mathbf{0}$ and $q_{2}=\mathbf{0}$ throughout training. Combined with a large-$N$ approximation for the empirical covariance $\hat{\Gamma}^{\mu}$ (which neglects the query’s contribution, see Appendix A.8.3, Eq. (A11)), the predictor for task $\mu$ takes a remarkably simpler form (derived in Appendix A.10, Eq. (A55)):
@@
Figure 2: Parameter dynamics: theory vs. simulation. (A) Empirical, batch-avergaed, evolution of the diagonal elements of the learned parameter $\bar{a}(t)=\bar{p}_{2}(t)\bar{q}_{1}(t)$ compared with theoretical predictions (Eq. 13), showing excellent agreement. (B) Mirror evolution of two terms in conserved quantity and confirmation of the stability of the conserved quantity $\mathcal{C}=\lVert\bar{p}_{2}\rVert_{F}^{2}-\lVert\bar{q}_{1}\rVert_{F}^{2}$.
@@
The system converges to a fixed point where $\frac{\mathrm{d}\bar{p}_{2}}{\mathrm{d}t}=\mathbf{0}$ and $\frac{\mathrm{d}\bar{q}_{1}}{\mathrm{d}t}=\mathbf{0}$. This occurs when the product of the parameters reaches $\bar{p}_{2}(\infty)\bar{q}_{1}(\infty)=[s^{\infty}(S)]^{-1}$. This fixed point result precisely matches those found in fixed-point analyses of linear transformers trained for in-context regression [8, 12]. Figure 1 confirms this theoretical prediction against simulations. The effective linear transformation learned by the model at convergence is derived by substituting this fixed point into the predictor (Eq. 8):
@@
In the large context limit ($N\to\infty$), assuming $\hat{\Sigma}_{x}\to\Sigma_{x}$, the term in brackets approaches $I$, yielding $\hat{y}_{\mu}(\infty)\to W^{\mu}x_{q}$. This demonstrates that the model learns to perfectly implement the target linear map. The term in brackets acts as a form of data-dependent preconditioning, related to ridge regression or preconditioned gradient descent updates identified in other works [8, 3].
Conserved Quantity and Reduced Degrees of Freedom. The ODE system (Eqs. 9-10) exhibits a conserved quantity: $\mathcal{C}=\lVert\bar{p}_{2}\rVert_{F}^{2}-\lVert\bar{q}_{1}\rVert_{F}^{2}$ remains constant throughout training (Figure 2B, Appendix A.12). This conservation law stems from an underlying scaling symmetry in the loss function and implies that the parameters evolve on a lower-dimensional manifold within the full parameter space, constraining their trajectory. This suggests a form of implicit regularization or low-dimensional learning, reminiscent of phenomena observed in other deep learning contexts [23, 24]. The existence of this conserved quantity can also be derived in a more principled way by observing that the loss function $\mathcal{L}\left(\bar{p}_{2},\bar{q}_{1}\right)$ is invariant under $\bar{p}_{2}\to\eta\bar{p}_{2}$ and $\bar{q}_{1}\to\frac{\bar{q}_{1}}{\eta}$, and using Noether’s theorem.
@@
where $a_{\alpha}^{\infty}=1/s_{\alpha}^{\infty}$ is the stable fixed point for mode $\alpha$, and $a_{\alpha}^{0}$ is the initial value. The characteristic time $t_{\alpha}^{*}$ required to transition from a small initialization $\epsilon$ to near its fixed point is approximately $t_{\alpha}^{*}\approx\frac{1}{2\lambda Ps_{\alpha}^{2}}\log\left(\frac{1}{s_{\alpha}^{\infty}\epsilon}\right)$. This inverse dependence on $s_{\alpha}^{2}$ establishes timescale separation: modes corresponding to larger input eigenvalues $s_{\alpha}$ learn significantly faster than those associated with smaller eigenvalues, mirroring the multi-stage learning observed in deep linear networks [10].
@@
This analytical form precisely captures the shape of the loss curve observed in simulations. The timescale separation inherent in $a_{\alpha}(t)$ directly translates into structured loss dynamics. The curve often exhibits plateaus, corresponding to periods where faster modes have converged but slower modes are still near their initial values (or slowly escaping saddle points near zero). Subsequent rapid drops ("cliffs") occur when these slower modes eventually learn, significantly reducing the remaining error (Figure 3A). This phenomenology of transformer training dynamics is widely observed empirically, and we show a particular example from a “grokking” Transformer [25] in Figure 3B, that has qualitative similarity to loss dynamics in our simplified model. The initial loss decreases linearly, with a rate proportional to initial parameter values and input covariance eigenvalues. The converged loss at $t\to\infty$ is non-zero due to the inherent randomness of finite context window and is given by $\mathcal{L}(\infty)=\frac{1}{2}\sum_{\alpha}\frac{s_{\alpha}(s_{\alpha}+\operatorname{Tr}(S))}{(N+1)s_{\alpha}+\operatorname{Tr}(S)}$. The loss dynamics near convergence are dominated by exponential relaxation, governed by the slowest learning modes. The final converged loss $\mathcal{L}(\infty)$ remains non-zero due to finite-$N$ effects related to noise in the empirical covariance estimates (Appendix A.13, Eq. (A114)).

### v2

@@ Abstract
keywords:
In-context learning, Linear Transformers, Learning Dynamics, Timescale Separation, Grokking
editors: List of editors’ names
@@ 1 Introduction
The transformer architecture (vaswani2017attention) has revolutionized machine learning, particularly through its capability for in-context learning (ICL). ICL allows models to adapt their behavior based on examples provided within the input context, obviating the need for task-specific parameter updates (brown2020language). This phenomenon links conceptually to meta-learning, characterized by distinct timescales: slow learning via weight updates during training and fast adaptation based on context during inference (vonOswald2023transformers). Unlike traditional supervised learning, ICL enables generalization to novel tasks by conditioning on contextual examples (chen2021evaluating). Mechanistically, attention layers can implement adaptive processing analogous to fast weights (schmidhuber1992learning; schlag2021linear; vonOswald2023transformers), suggesting that transformers might internally simulate learning algorithms (akyurek2023context; vonOswald2023transformers; ahn2023transformers). Despite compelling empirical demonstrations, a full theoretical grasp of how transformers achieve ICL remains unclear.
To address this challenge, our work presents an in-depth analysis of a simplified, mathematically tractable transformer model performing in-context linear regression. Our model incorporates specific simplifications—linear attention (vonOswald2023transformers; akyurek2023context), a structured input embedding (garg2022can; akyurek2023context), mean-squared error loss, and a continuous-time approximation of learning dynamics—yet retains core aspects of ICL. This deliberate simplification allows us to analytically trace how contextual information modulates the model’s predictions. By leveraging methods from the study of exactly solvable models, introduced first in deep linear networks (saxe2014exact), we derive a precise characterization of the learning dynamics under stochastic gradient descent (SGD). We analyze the resulting phenomenology, including the emergence of multiple learning timescales dictated by data statistics (saxe2014exact; rahaman2019spectral). This detailed understanding of a simplified system aims to provide basic insights applicable to the interpretation and analysis of more complex transformer models exhibiting ICL. In particular, we believe that this detailed understanding can inform interpretability efforts and evaluation methodologies for more complex, real-world models and argue the point by quantitatively analyzing the properties of model parameters in 1) simple attention only transformer model that exhibit sudden emergence of in context learning during training time, and 2) grokking model that exhibits delayed generalization.
@@ 1.1 Related Work
Transformers exhibit remarkable in-context learning (ICL), adapting to new tasks using context examples without explicit parameter updates (brown2020language). This connects to meta-learning’s dual timescales: slow weight learning and fast context adaptation (vonOswald2023transformers). Early concepts like fast weights (schmidhuber1992learning) suggested context-based adaptation mechanisms. Recent studies demonstrate that linear self-attention layers can be mathematically equivalent to fast weight update rules (schlag2021linear; vonOswald2023transformers), enabling Transformers to implement learning algorithms within their forward pass (vonOswald2023transformers).
@@
Several studies confirm that Transformers can encode standard learning algorithms. akyurek2023context proved by construction that Transformers can implement gradient descent (GD) and ridge regression, finding empirically that trained models mimic these algorithms and converge towards Bayesian optimal predictors for linear tasks. Similarly, vonOswald2023transformers demonstrated that Transformers trained on regression problems learn GD internally, with attention updates mapping directly to GD steps, effectively becoming ”mesa-optimizers”. In parallel work ahn2023transformers, and zhang provided theoretical guarantees that standard training yields preconditioned GD at the fixed-point of the training dynamics across a range of setups. These findings collectively suggest that standard training procedures can induce Transformers to implicitly learn optimization algorithms. Extending this, fu2024transformers found evidence that Transformers might learn higher-order methods resembling Newton’s method, explaining superior performance on ill-conditioned problems. Alternative perspectives conceptualize ICL as Bayesian inference, where models infer latent task variables from prompts (xie2022explanation), or as kernel regression, where attention patterns reflect similarity-based weighting akin to kernel methods (han2023explaining). These views provide explanations for observed phenomena like prompt sensitivity and the utility of relevant examples (xie2022explanation; han2023explaining). The precise mechanism—gradient-based, Bayesian, or kernel-like—likely depends on model scale, architecture, and training data (akyurek2023context; vonOswald2023transformers; han2023explaining; xie2022explanation; ahn2023transformers; zhang).
@@
At the same time, analyzing learning dynamics has a rich history, notably through the study of deep linear networks (saxe2014exact). saxe2014exact derived exact solutions for these models, revealing nonlinear dynamics and multiple timescales despite linearity. Learning often occurs in stages, marked by plateaus followed by rapid error reduction as different eigenmodes of the data correlation are sequentially learned. Learning timescales are typically eigenvalue-dependent, with dominant modes learned faster (saxe2014exact; advani2017high). In nonlinear networks, exact solutions are rare, but phenomena like spectral low rank bias indicate structured learning dynamics: models tend to learn low-frequency (smooth) function components before high-frequency details (rahaman2019spectral). This bias is linked to the Neural Tangent Kernel (NTK) spectrum, where larger eigenvalues associated with smoother functions lead to faster learning (rahaman2019spectral; Jacot2018Neural), potentially explaining the generalization capabilities of overparameterized models.
@@
Separation of timescales emerges as a common theme across linear dynamics (saxe2014exact), nonlinear spectral bias (rahaman2019spectral), meta-learning (vonOswald2023transformers) and the abrupt emergence of ICL or induction heads (olsson2022context) hints at similar timescale separation. Indeed, there have been theoretical results in non-linear transformer establishing distinct phases in the learning dynamics boix-adsera; chen2024 further supporting the idea. Alternative models like associative memories instead connect ICL to retrieval dynamics (ramsauer2021hopfield).
Most recently, parallel studies have also explored the theoretical properties of linear attention. lu2025 and lyu2025 similarly identify that linear attention models learn to whiten or precondition data; however, these works focus primarily on asymptotic behavior and scaling laws in multi-task regimes, respectively. Additionally, zhang2025 analyze the exact learning dynamics for scalar outputs ($\mathbb{R}^{d}\to\mathbb{R}^{1}$) within multi-head architectures and different QK parametrization. In contrast, our work provides an non-asymptotic closed-form characterization of the learning dynamics for vector-valued ($\mathbb{R}^{d}\to\mathbb{R}^{d}$) regression and apply the insights from the theory for interpreting non-linear models.
@@ 2.1 Linear Transformer Architecture
We analyze a single-layer transformer with a single linear attention head. In contrast to standard architectures using softmax attention (vaswani2017attention), we employ simplifications common in theoretical studies (vonOswald2023transformers; akyurek2023context). The model utilizes two weight matrices, $W^{Q}\in\mathbb{R}^{D\times D}$ (query/key) and $W^{P}\in\mathbb{R}^{D\times D}$ (output/projection), and replaces the attention non-linearity with the identity function ($\phi(x)=x$). The layer’s computation is defined as (see Appendix A.1-A.3, Eq. (A3)):
@@
Here, $Z\in\mathbb{R}^{D\times(N+1)}$ is the input sequence embedding, and $N+1$ is the sequence length. The learnable parameters are $\Theta=\{W^{P},W^{Q}\}$. This linear attention model, while simplified, is sufficient for learning in-context linear regression (vonOswald2023transformers).
Relevant parameter submatrices are defined based on the input embedding structure (Appendix A.2). We partition $W^{Q}$ such that its first $d$ columns form $q=\left[q_{1}^{T}\,\mid\,q_{2}^{T}\right]^{T}$, and $W^{P}$ such that its bottom $d$ rows form $p^{T}=\left[p_{1}\,\mid\,p_{2}\right]$. All submatrices $q_{1},q_{2},p_{1},p_{2}$ are in $\mathbb{R}^{d\times d}$. More precisely, we have the following parametrization:
$W^{P}$ $=\begin{pmatrix}\cdot&\cdot\\ p_{1}&p_{2}\end{pmatrix}\quad\text{with}\quad p^{T}=\begin{pmatrix}p_{1}&p_{2}\end{pmatrix}$, (2)
@@ 2.2 Embedding Structure and Task Setup for In-Context Regression
The model is trained to perform in-context linear regression. Given a prompt containing $N$ example pairs $(x_{i},y_{i})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, it must predict the output $y_{q}$ for a query input $x_{q}\in\mathbb{R}^{d}$ based on the linear relationship demonstrated in the prompt. This setup extends the $\mathbb{R}^{d}\to\mathbb{R}^{1}$ tasks studied in some prior work (akyurek2023context; garg2022can) to the more general vector-valued regression setting $\mathbb{R}^{d}\to\mathbb{R}^{d}$.
We employ a structured embedding strategy commonly used in ICL studies (akyurek2023context; garg2022can; raffel2020exploring). For an input sequence $(x_{1},y_{1},\dots,x_{N},y_{N},x_{q})$, the embedding matrix $Z\in\mathbb{R}^{D\times(N+1)}$ (with $D=2d$) is constructed as:
@@ 3 Theory
Simplified predictor. The model predicts the query output $\hat{y}\in\mathbb{R}^{d}$ using the bottom $d$ elements of the final output column $f(Z)]_{:,N+1}$. As derived in Appendix A.9 (Eq. (A18)), this prediction simplifies to a linear transformation of the query input:
@@
Null Parameter Dynamics and Predictor Simplification. A key simplification arises from identifying $p_{1}$ and $q_{2}$ as ”null parameters.” Unlike some prior theoretical works which simplify the optimization landscape by assuming specific parameterizations (e.g., rank constraints or symmetric weight assumptions) a priori, our simplification is derived dynamically. We show in Appendix A.10 that if initialized at zero, the expected gradients for $p_{1}$ and $q_{2}$ vanish identically zhang; ahn2023transformers. This implies that the dynamics naturally confine the optimization to the relevant subspace without enforcing hard constraints, theoretically justifying fixing $p_{1}=\mathbf{0}$ and $q_{2}=\mathbf{0}$. Combined with a large-$N$ approximation for the empirical covariance $\hat{\Gamma}^{\mu}$ (which neglects the query’s contribution, see Appendix A.8.3, Eq. (A11)), the predictor for task $\mu$ takes a remarkably simpler form (derived in Appendix A.10, Eq. (A55)):
@@
Figure 2: Parameter dynamics: theory vs. simulation. (A) Empirical, batch-averaged, evolution of the diagonal elements of the learned parameter $\bar{a}(t)=\bar{p}_{2}(t)\bar{q}_{1}(t)$ compared with theoretical predictions (Eq. 13), showing excellent agreement. (B) Mirror evolution of two terms in conserved quantity and confirmation of the stability of the conserved quantity $\mathcal{C}=\lVert\bar{p}_{2}\rVert_{F}^{2}-\lVert\bar{q}_{1}\rVert_{F}^{2}$.
@@
The system converges to a fixed point where $\frac{\mathrm{d}\bar{p}_{2}}{\mathrm{d}t}=\mathbf{0}$ and $\frac{\mathrm{d}\bar{q}_{1}}{\mathrm{d}t}=\mathbf{0}$. This occurs when the product of the parameters reaches $\bar{p}_{2}(\infty)\bar{q}_{1}(\infty)=[s^{\infty}(S)]^{-1}$. This fixed point result precisely matches those found in fixed-point analyses of linear transformers trained for in-context regression (ahn2023transformers; zhang). Figure 1 confirms this theoretical prediction against simulations. The effective linear transformation learned by the model at convergence is derived by substituting this fixed point into the predictor (Eq. 8):
@@
In the large context limit ($N\to\infty$), assuming $\hat{\Sigma}_{x}\to\Sigma_{x}$, the term in brackets approaches $I$, yielding $\hat{y}_{\mu}(\infty)\to W^{\mu}x_{q}$. This demonstrates that the model learns to perfectly implement the target linear map. The term in brackets acts as a form of data-dependent preconditioning, related to ridge regression or preconditioned gradient descent updates identified in other works (ahn2023transformers; vonOswald2023transformers).
Conserved Quantity and Reduced Degrees of Freedom. The ODE system (Eqs. 9-10) exhibits a conserved quantity: $\mathcal{C}=\lVert\bar{p}_{2}\rVert_{F}^{2}-\lVert\bar{q}_{1}\rVert_{F}^{2}$ remains constant throughout training (Figure 2B, Appendix A.12). This conservation law stems from an underlying scaling symmetry in the loss function and implies that the parameters evolve on a lower-dimensional manifold within the full parameter space, constraining their trajectory. This suggests a form of implicit regularization or low-dimensional learning, reminiscent of phenomena observed in other deep learning contexts (biroli2022kernel; gerace2020disentangling). The existence of this conserved quantity can also be derived in a more principled way by observing that the loss function $\mathcal{L}\left(\bar{p}_{2},\bar{q}_{1}\right)$ is invariant under $\bar{p}_{2}\to\eta\bar{p}_{2}$ and $\bar{q}_{1}\to\frac{\bar{q}_{1}}{\eta}$, and using Noether’s theorem.
@@
where $a_{\alpha}^{\infty}=1/s_{\alpha}^{\infty}$ is the stable fixed point for mode $\alpha$, and $a_{\alpha}^{0}$ is the initial value. The characteristic time $t_{\alpha}^{*}$ required to transition from a small initialization $\epsilon$ to near its fixed point is approximately $t_{\alpha}^{*}\approx\frac{1}{2\lambda Ps_{\alpha}^{2}}\log\left(\frac{1}{s_{\alpha}^{\infty}\epsilon}\right)$. This inverse dependence on $s_{\alpha}^{2}$ establishes timescale separation: modes corresponding to larger input eigenvalues $s_{\alpha}$ learn significantly faster than those associated with smaller eigenvalues, mirroring the multi-stage learning observed in deep linear networks (saxe2014exact).
@@
This analytical form precisely captures the shape of the loss curve observed in simulations. The timescale separation inherent in $a_{\alpha}(t)$ directly translates into structured loss dynamics. The curve often exhibits plateaus, corresponding to periods where faster modes have converged but slower modes are still near their initial values (or slowly escaping saddle points near zero). Subsequent rapid drops (”cliffs”) occur when these slower modes eventually learn, significantly reducing the remaining error (Figure 3A). This phenomenology of transformer training dynamics is widely observed empirically, and we show a particular example from a “grokking” Transformer (nanda2023grokking) in Figure 3B, that has qualitative similarity to loss dynamics in our simplified model. The initial loss decreases linearly, with a rate proportional to initial parameter values and input covariance eigenvalues. The converged loss at $t\to\infty$ is non-zero due to the inherent randomness of finite context window and is given by $\mathcal{L}(\infty)=\frac{1}{2}\sum_{\alpha}\frac{s_{\alpha}(s_{\alpha}+\operatorname{Tr}(S))}{(N+1)s_{\alpha}+\operatorname{Tr}(S)}$. The loss dynamics near convergence are dominated by exponential relaxation, governed by the slowest learning modes. The final converged loss $\mathcal{L}(\infty)$ remains non-zero due to finite-$N$ effects related to noise in the empirical covariance estimates (Appendix A.13, Eq. (A114)).